# Unlocking Korean Verbs: A User-Friendly Exploration into the Verb Lexicon

### v2

@@ A User-Friendly Exploration into the Verb Lexicon
Seohyun Song1∗ Eunkyul Leah Jo2∗ Yige Chen3∗ Jeen-Pyo Hong4††thanks: Equally contributed authors. Corresponding authors: ktlim@seoultech.ac.kr and jungyeul@mail.ubc.ca. Kyuwon Kim5 Jin Wee6 Miyoung Kang6 KyungTae Lim1† Jungyeul Park2† Chulwoo Park7 1SeoulTech, South Korea 2The University of British Columbia, Canada 3The Chinese University of Hong Kong, Hong Kong 442dot Inc., South Korea 5Seoul National University, South Korea 6National Institute of Korean Language, South Korea 7Anyang University, South Korea https://www.korean.go.kr
@@ 1 Introduction
When examining the verb lexicon for Korean, the Sejong dictionary has been a major language resource. As part of the 21st Century Sejong Project, which aims to advance Korean language information processing, the Sejong dictionary has produced extensive datasets that describe Korean lexicon data in great detail. Although it is designed with a vast amount of morphological, semantic and syntactic information, there have been minimal attempts to utilize its resources. This paper investigates the Korean verb lexicon, with a particular focus on the subcategorization frames in the Sejong verb dictionary. The objective is to facilitate easier access to and manipulation of the Sejong dictionary, thereby promoting further studies and applications in Korean language processing.
@@
Within the Sejong dictionary are XML files for each verb (Figure 1). Inside are XML tags that identify each word’s morphological, semantic, and syntactic information with extra metadata. Some annotations make it possible to identify arguments such as semantic roles of verbs in sentences (Figure 2). A set of the underlying semantic roles for each verb and frame groupings are defined for each sense of the verb. Then, in each group are the frames that annotate selection restrictions with example sentences. We believe this detailed verb representation level is vital for many language processing applications. With thorough and manual documentation on subcategorization in the Sejong dictionary, we aim to enhance the dataset’s usability, facilitating the creation of user-friendly systems for all future projects.
@@ 2.2 Previous usages of the Sejong dictionary
The Sejong dictionary has been adopted in a limited way. Data and example sentences from the Sejong dictionary were used in the Korean SRL dataset developed by the National Institute of Korean Language, which was later improved through a linguistically-informed annotation strategy (Chen et al., 2024b). Additionally, a small selection of example sentences from the Sejong dictionary was incorporated into Korean FrameNet (Hahm et al., 2018; Chen et al., 2024a), guided by the frames outlined in the dictionary entries within the Sejong dictionary.222https://github.com/machinereading/koreanframenet,333https://github.com/jungyeul/k-framenet
@@ 3 Developing a Web Interface
The argument boundaries are detected through a combination of a heuristic-based chunking method and the sentence’s dependency structures. Argument boundaries are determined by the locations of postpositions in the sentence, as postpositions in Korean typically indicate the endpoints of arguments and modifiers, based on our previous work (Chen et al., 2024b). On the other hand, dependency structures assist in identifying argument boundaries when the heuristics are insufficient to derive chunks for certain constituents (Chen et al., 2022). The postpositions of the chunks are further compared with the postpositions given by the frame to determine whether a chunk is an argument or not.
@@ 5 Conclusion and Future Perspectives
Our current contributions, particularly the Python library, have the potential to be applied across various linguistic tasks, including syntactic parsing systems that utilize lexical information, such as lexicalized parsers (Sarkar and Han, 2002; Park et al., 2013), semantic role labeling (Changki et al., 2015; Seo-In, 2019; Chen et al., 2024b), and FrameNet parsing (Park et al., 2014; Kim et al., 2016; Hahm et al., 2018; Chen et al., 2024a) for the Korean language. We are committed to broadening access to Korean language data beyond traditional boundaries, offering a unified platform for linguists and computer scientists to explore the rich resources of the Korean language. This platform will be invaluable for a wide range of Korean language processing tasks and future linguistic research.
We are continuously working on integrating additional Korean verb lexicons, such as the Korean PropBank and FrameNet, into our web service. Our ultimate goal is to develop a comprehensive Korean VerbNet. By systematically comparing verbs and their subcategorization frames across the Sejong verb dictionary, PropBank, and other resources—and resolving any discrepancies—we aim to harmonize frame information across our interfaces. This will offer users a more cohesive and seamless experience when exploring Korean language data, contributing to advancements in Korean linguistic research and natural language processing.

### v3

@@ A User-Friendly Exploration into the Verb Lexicon
Seohyun Song1∗ Eunkyul Leah Jo2∗ Yige Chen3∗ Jeen-Pyo Hong4††thanks: Equally contributed authors. Corresponding authors: KyungTae Lim and Jungyeul Park. Kyuwon Kim5 Jin Wee6 Miyoung Kang6 KyungTae Lim7† Jungyeul Park2† Chulwoo Park8 1Seoultech, South Korea 2The University of British Columbia, Canada 3The Chinese University of Hong Kong, Hong Kong 442dot Inc., South Korea 5Seoul National University, South Korea 6National Institute of Korean Language, South Korea 7KAIST, South Korea 8Anyang University, South Korea https://www.korean.go.kr
@@ 1 Introduction
When examining the verb lexicon for Korean, the Sejong dictionary has been a major language resource. As part of the 21st Century Sejong Project, which aims to advance Korean language information processing, the Sejong dictionary has produced extensive datasets that describe Korean lexicon data in great detail. Although it is designed with a vast amount of morphological, semantic, and syntactic information, there have been minimal attempts to utilize its resources. This paper investigates the Korean verb lexicon, with a particular focus on the subcategorization frames in the Sejong verb dictionary. The objective is to facilitate easier access to and manipulation of the Sejong dictionary, thus promoting further studies and applications in Korean language processing.
@@
Within the Sejong dictionary are XML files for each verb (Figure 1). Inside are XML tags that identify each word’s morphological, semantic, and syntactic information with additional metadata. Some annotations make it possible to identify arguments such as semantic roles of verbs in sentences (Figure 2). A set of the underlying semantic roles for each verb and frame groupings are defined for each sense of the verb. Then, in each group are the frames that annotate selection restrictions with example sentences. We believe this detailed verb representation level is vital for many language processing applications. With thorough and manual documentation on subcategorization in the Sejong dictionary, we aim to enhance the dataset’s usability, facilitating the creation of user-friendly systems for all future projects.
@@ 2.2 Previous usages of the Sejong dictionary
The Sejong dictionary has been adopted in a limited way. Data and example sentences from the Sejong dictionary were used in the Korean SRL dataset developed by the National Institute of Korean Language, which was later improved through a linguistically-informed annotation strategy (Chen et al., 2024c). Additionally, a small selection of example sentences from the Sejong dictionary was incorporated into Korean FrameNet (Hahm et al., 2018; Chen et al., 2024a), guided by the frames outlined in the dictionary entries within the Sejong dictionary.222https://github.com/machinereading/koreanframenet,333https://github.com/jungyeul/k-framenet
@@ 3 Developing a Web Interface
The argument boundaries are detected through a combination of a heuristic-based chunking method and the sentence’s dependency structures. Argument boundaries are determined by the locations of postpositions in the sentence, as postpositions in Korean typically indicate the endpoints of arguments and modifiers, based on our previous work (Chen et al., 2024c). On the other hand, dependency structures assist in identifying argument boundaries when the heuristics are insufficient to derive chunks for certain constituents (Chen et al., 2022). The postpositions of the chunks are further compared with the postpositions given by the frame to determine whether a chunk is an argument or not.
@@ 5 Conclusion and Future Perspectives
Our current contributions, particularly the Python library, have the potential to be applied across various linguistic tasks, including syntactic parsing systems that utilize lexical information, such as lexicalized parsers (Sarkar and Han, 2002; Park et al., 2013), semantic role labeling (Changki et al., 2015; Seo-In, 2019; Chen et al., 2024c, b), and FrameNet parsing (Park et al., 2014; Kim et al., 2016; Hahm et al., 2018; Chen et al., 2024a) for the Korean language. We are committed to broadening access to Korean language data beyond traditional boundaries, offering a unified platform for linguists and computer scientists to explore the rich resources of the Korean language. This platform will be invaluable for a wide range of Korean language processing tasks and future linguistic research.
We are continuously working on integrating additional Korean verb lexicons, such as the Korean PropBank and FrameNet, into our web service. Our ultimate goal is to develop a comprehensive Korean VerbNet. By systematically comparing verbs and their subcategorization frames across the Sejong verb dictionary, PropBank, and other resources—and resolving any discrepancies—we aim to harmonize frame information across our interfaces. This will offer users a more cohesive and seamless experience when exploring Korean language data, contributing to advancements in Korean linguistic research and natural language processing. Our system, including its Python library with sample verb files, can be accessed at http://117.17.185.27:8025/.